13.9grams of an unknown gas is placed in a 5.00L container. It has an initial pressure at 58.6kPa and initial temperature at 60.0°C. What is the name of this gas?
1
Expert's answer
2020-10-19T14:02:27-0400

PV = nRT – ideal gas law

P = 58.6 kPa

V = 5 L

T = 60 ºC + 273.15 = 333.15 K

R =8.31447 L kPa mol-1 K-1

"n = \frac{PV}{RT}"

"n = \frac{58.6 \times 5}{8.31447 \times 333.15} = 0.1057\; mol"

M = \frac{m}{n}

m = 13.9 g

"M(gas) = \frac{13.9}{0.1057} = 131.5\; g/mol"

Examining a periodic table leads us to identifying the gas as xenon.

Answer: xenon.